# Active Matter as a framework for living systems-inspired Robophysics

### v1

@@ I Introduction
The development of robots has been a growing focus of the scientific community for over half a century, with particularly rapid advances in the past three decades Aguilar et al. (2016); Wang et al. (2021). These versatile, remotely controlled, or autonomous devices are increasingly essential due to their ability to perform tasks that are dangerous, repetitive, or difficult for humans. Robots offer significant benefits in tasks such as search and rescue operations Lyu et al. (2023), being able to access unsafe environments, as well as damage evaluation, agricultural monitoring, forest management Ecke et al. (2022), and crowd monitoring Husman et al. (2021); Conte et al. (2021); Al-Sheary and Almagbile (2017); Gohari et al. (2022), enhancing precision and efficiency. Despite their remarkable capabilities and diverse applications, robots still encounter substantial efficiency limitations both as individual units and in collective operations. In comparison to biological organisms, at the single-robot level, locomotion remains a critical challenge: robots often struggle to navigate natural terrains, such as dense forests, rubble-strewn areas, or regions with gusty windsAguilar et al. (2016); Schaeffer and Lindstedt (2013); Fajen (2013); Biewener and Patek (2018). Additionally, mission planning is frequently challenging due to the robot’s inability to effectively manage unforeseen obstacles, resulting in decreased operational efficiency and reliability. To address these challenges, recent research efforts have focused on incorporating principles from physics and biology into the modeling and design of robotsVolpe et al. (2025).
@@
The gap in understanding motion within complex environments, particularly from a physics perspective, has given rise to the field of Robophysics Aguilar et al. (2016). Robophysics is defined as the pursuit of principles governing the movement and control of robot-like entities in intricate environments. It delves into fundamentally non-equilibrium processes where motion emerges through internal actuation Levine and Goldman (2023). Advancing the understanding of robotic movement principles necessitates the application of experimental physics, combining theoretical and numerical frameworks with systematically controlled laboratory experiments on simplified systems.
@@
In this perspective article, we examine how Robophysics provides a framework for understanding and improving both individual and collective robotic behaviors. We first discuss how insights from living systems have informed the design of single-robot locomotion strategies. We then explore the extension of these principles to collective systems, highlighting the challenges that arise in robot swarms. Finally, we outline how recent advances in active matter physics and machine learning can be integrated to develop predictive, adaptive, and purpose-driven robotic collectives.
II Learning from living systems: single robot
@@
At the single-robot level, Robophysics has focused on understanding and improving locomotion by building simplified devices capturing essential geometric and control features of biological systems Aguilar et al. (2016). These devices function as abstract models, providing insights into appendage-induced locomotion, blurring the dichotomy between failure and success to uncover principles of environment-locomotor interaction Aguilar et al. (2016). Consequently, the exploration of robotic motion serves as an exemplary model system for advancing our understanding of complex dynamical systems driven far from equilibrium Bechinger et al. (2016); Volpe et al. (2025). So far, the scientific literature has reported both terrestrial and aerial examples in which Robophysics played a pivotal role in enabling robots to adapt to real-world conditions by incorporating living-like mechanisms into their design.
@@
On the one side, on land, snake-like robots have been designed to move through unpredictable and confined spaces by mimicking the undulating motion of snakes, which use their body segments and frictional forces to generate propulsion Takemori et al. (2018); Tanaka et al. (2015). This ability makes snake-like robots highly suitable for deployment in disaster zones, where they can traverse rubble and debris to locate survivors or assess structural damageQiyuan and Chen (2020); Fu et al. (2020). Studies of snake locomotion across diverse terrains have allowed the creation of robots excelling in mobility, adaptability, and durability, even under extreme conditions Moosavi et al. (2025); Thanabalan et al. (2025). At a smaller length-scale, peristaltic warm-inspired soft robots have been designed for a better understanding of locomotion in different environments Das et al. (2023).
@@
On the other side, in the air, bird-inspired drones have been tailored to exploit flapping-wing aerodynamics to achieve maneuverability and energy efficiency which surpasses traditional fixed-wing or rotor designsPhan and Floreano (2024). At a smaller length scale, insect-inspired robots have been designed for crop pollination and environmental monitoringChen et al. (2019). Such bio-inspired abstractions demonstrate how Robophysics, by bridging biology and engineering, enhances robotic adaptability in complex real-world environments Jeger et al. (2024).
@@
III Learning from living systems: collective behaviour
@@
Living systems achieve collective functionality not via a central control, but via local interactions which generate robust and adaptable group behavior Hauser (1997); Gillam (2011). Across length-scales, organisms actively consume energy to remain out of equilibrium and thereby dynamically respond to environmental cues Wang et al. (2015). For example, cytoskeletal filaments continuously assemble and disassemble to enable dynamic cell shape changes Fletcher and Mullins (2010), while flocks of birds or schools of fish reorganize within fractions of a second to conserve energy or evade predators Marchetti et al. (2013); Couzin et al. (2005); Portugal et al. (2014). The unifying principle is that coordinated responses emerge from simple, local rules of sensing and interaction, yet giving rise to collective resilience and adaptability.
@@
Efficient robot swarms rely on fundamental processes-communication, coordination, cooperation, collaboration, and, at times, competition Prorok et al. (2021), to achieve collective goals. Algorithms based on these tasks have already achieved remarkable results in controlled experimental environments with a limited number of individuals in the swarm (fewer than ten). The following sections discuss each of these aspects and their relevance to collective robotic behavior.
III.1 Communication and its constraints
@@
Gielis et al. Gielis (2022) argue that communication in robot swarms confronts two major challenges. First, learning models should be employed not only to determine what information to communicate and to whom, but also to optimize multiple aspects of swarm performance simultaneously. Second, robots must be designed together with their communication capabilities, since any control algorithm relying on explicit communication is prone to failure when network data are insufficient and to sub-optimal performance when the network transmits excessive information.
At a first glance, Robophysics may appear similar to bio-inspired networking, as both draw on concepts from collective behavior in natural systems. However, bio-inspired networking primarily aims to leverage biological models to improve the performance of routing, coordination, synchronization, and resource-allocation algorithms in large-scale communication networks Dressler and Akan (2010). Robophysics, instead, seeks to develop a physics-based approach of how real robots move, interact, and generate collective behaviors under physical constraints. From a computer networking perspective, the operation of a swarm of small, low-cost robots faces several fundamental challenges. Unlike abstract network nodes, robots are physical entities whose positions and motions directly affect network connectivity. The network topology continuously evolves with the motion of the swarm, requiring tightly coupled network strategies for locomotion and control.
However, a location-deprived environment, as is typical in many indoor robophysical experiments, poses an additional challenge to robots in terms of simultaneous location and mapping (SLAM), which is necessary for each robot to estimate both its relative position and the overall network topology. Typical SLAM systems assume relatively high-cost, high-performance hardware (e.g., LiDAR) and often cannot be deployed effectively on ultra-low-cost robotic platforms due to high power, mechanical complexity, and computational demands.
Performing topology mapping and cooperative localization in indoor environments with only simple sensors is thus a significant challenge. Relying on low-cost sensors inherently introduces sensor uncertainty, since measurements of distance, velocity, or orientation are often noisy or incomplete. This, in turn, leads to perception uncertainty, as obstacles or neighboring robots may be only partially detected or misidentified due to occlusions or reflections. Moreover, robot mobility exacerbates communication uncertainty, as time-varying and intermittent wireless links can cause packet losses and transmission delays.
In Robophysics, however, communication is not merely an abstract tool but an integral component of the physical process itself: information exchange influences motion, and motion, in turn, shapes connectivity. Modeling and controlling this bidirectional feedback loop remains an open challenge situated at the intersection of network dynamics, control theory, and collective behavior.
III.2 Coordination and Scalability
Beyond communication, an equally critical challenge lies in coordination, which seeks to minimize redundant tasks and interference through cooperative behaviors such as path synchronization for area coverage and collision avoidance Yao et al. (2022).
@@ III.2 Coordination and Scalability
Designing a robot swarm requires scalability as the number of robots increases. Yet, many techniques that perform well for small groups become impractical when the swarm size grows. Recent research has therefore focused on the formal control of shape formation in robot collectives Sakurama and Sugie (2021). In these systems, robots communicate with their local neighbors to maintain a desired configuration, and their connections are typically represented as a graph. Scaling up such formations involves enlarging the graph while preserving its topological properties, a task that remains particularly challenging.
@@
Figure 1: Schematic illustration of how principles from living systems can inform robophysics. In living systems (top), groups of organisms such as fish, ants, and birds display collective behavior, adaptability, and shared purpose. These principles inspire robotic collectives (bottom), where multi-robot swarms must address challenges such as communication, perception, and sensing, and energy and resource management.
III.3 Cooperation, routing and planning
@@
Cooperation refers to the ability of similar agents to work together on tasks that cannot be accomplished individually, such as jointly moving a payload or assisting one another in crossing a gap. For many years, the field of computer networking has drawn ideas from bio-inspired models of collective behavior, leading to the development of algorithmic paradigms such as ant-colony-inspired routing and optimization schemes. In ant-based routing protocols for mobile ad hoc networks, artificial ’ants’ (data packets) explore paths and probabilistically reinforce successful ones, mimicking pheromone trail deposition. This approach, known as Ant Colony Optimization (ACO), has proven effective in dynamic, decentralized settings by leveraging self-organization and adaptability to changing network topologies.
@@
Similarly, evolutionary algorithms such as genetic algorithms (GAs) have been applied to path planning and task assignment in multi-drone systems. For example, GA-based strategies like GenPath Bartolini et al. (2021) optimize UAV (Unmanned Aerial Vehicle) missions under constraints such as limited battery life, intermittent communication, and multi-phase recharging, enabling scalable coordination in search-and-rescue, environmental monitoring, and parcel delivery.
While ant colony optimization and genetic algorithms excelled at solving static or well-defined problems of routing and planning, they did not capture the real-time, embodied responsiveness of biological collectives, where coordination emerges from local physical interactions and rapid adaptation to changing conditions.
This gap highlights the relevance of Robophysics to adopt a different point of view: learning from living systems to program robots whose collective behavior emerges in analogous ways. According to this view, coordination is not imposed top-down but arises from the living-inspired physics of local interactions, fluctuations, and energy consumption.
III.4 Collaboration, Learning and Adaptation
@@
Collaboration emphasizes teamwork with a diverse team that leverages the unique skills of different individuals, such as integrating various sensors or carrying out a common patrol task combining aerial and aquatic autonomous vehicles.
Learning concerns the swarm’s ability to learn from new scenarios and adjust its collective behaviour to achieve its goals, which is crucial for effective operation. While learning, robots in a swarm can collectively perceive their surroundings and share that information among themselves (distributed perception and sensing).
@@
To better adapt to the environment, robots should exhibit both robustness and fault tolerance. Although these properties are related, they represent distinct concepts. Robustness refers to the ability to maintain satisfactory performance even when environmental conditions differ from those anticipated during the design phase. Fault tolerance, instead, refers to the capacity of the swarm to sustain effective operation despite failures in individual agents, sensors, or other components.
III.5 Competition and a shared purpose
@@
In both biological and robotic collectives, competition can coexist with cooperation, as individuals pursuing local goals may still contribute to global functionality. Such emergent coordination depends not only on feedback regulation but on a shared purpose. At the microscopic scale, gene regulatory networks couple local signals to coordinated multicellular responses during processes like morphogenesis, embodying distributed feedback Grada et al. (2017); Angelini et al. (2011); Bi et al. (2016). At the macroscopic scale, local interaction rules (e.g., alignment, attraction, repulsion) are goal-conditioned: animals modulate the way they weigh neighbors and environmental cues in service of collective functions such as predator evasion, energy saving, or foraging Couzin et al. (2005); Portugal et al. (2014); Reebs (2000); Swaney et al. (2001). Robustness in living systems thus arises from distributed decision-making with a purpose, real-time adaptability driven by local feedback tuned to a common objective, rather than pre-planned, centralized control Marshall and Reina (2024).
@@
III.6 Testing robot swarms

### v2

# Dept. Arquitectura de Computadores y Automatica, Facultad de Ciencias Fisicas, Universidad Complutense 28040 Madrid, Spain Departamento de Estructura de la Materia, Física Térmica y Electrónica, Universidad Complutense de Madrid, 28040 Madrid, Spain GISC - Grupo Interdisciplinar de Sistemas Complejos 28040 Madrid, Spain

@@ 1 Introduction
The development of robots has been a growing focus of the scientific community for over half a century, with particularly rapid advances in the past three decades [3, 89]. Pioneeringly, more than two decades ago, Brooks and Flynn proposed replacing traditional expensive planetary exploration with fleets of small, autonomous, low-cost robots [19]. Such versatile devices — remotely operated or fully autonomous — are becoming increasingly essential, due to their ability to perform tasks which are dangerous, repetitive, or otherwise challenging for humans. Robots offer significant benefits in tasks such as search and rescue operations [59], being able to access unsafe environments, as well as damage evaluation, agricultural monitoring, forest management [34], and crowd monitoring [48, 27, 4, 44], enhancing precision and efficiency. Despite their remarkable capabilities and diverse applications, robots still encounter substantial efficiency limitations both as individual units and in collective operations. In comparison to biological organisms, at the single-robot level, locomotion remains a critical challenge: robots often struggle to navigate natural terrains, such as dense forests, rubble-strewn areas, or regions with gusty winds[3, 79, 35, 14]. Additionally, mission planning is frequently challenging due to the robot’s inability to effectively manage unforeseen obstacles, resulting in decreased operational efficiency and reliability. To address these challenges, recent research efforts have focused on incorporating principles from physics and biology into the modeling and design of robots[88].
@@
The gap in understanding motion within complex environments, particularly from a physics perspective, has given rise to the field of Robophysics [3]. Robophysics is defined as the pursuit of principles governing the movement and control of robot-like entities in intricate environments. It delves into fundamentally non-equilibrium processes where motion emerges through internal actuation [55]. Advancing the understanding of robotic movement principles necessitates the application of experimental physics, combining theoretical and numerical frameworks with systematically controlled laboratory experiments on simplified systems.
@@
In this perspective article, we examine how Robophysics provides a framework for understanding and improving both individual and collective robotic behaviors. We first discuss how insights from living systems have informed the design of single-robot locomotion strategies. Next, we explore the extension of these principles to collective systems, highlighting the challenges that arise in robot swarms. Finally, we outline how recent advances in active matter physics and machine learning can be integrated to develop predictive, adaptive, and purpose-driven robotic collectives.
2 Learning from living systems: single robot
@@
At the single-robot level, Robophysics has focused on understanding and improving locomotion by building simplified devices capturing essential geometric and control features of biological systems [3]. These devices function as abstract models, providing insights into appendage-induced locomotion, blurring the dichotomy between failure and success to uncover principles of environment-locomotor interaction [3]. Consequently, the exploration of robotic motion serves as an exemplary model system for advancing our understanding of complex dynamical systems driven far from equilibrium [11, 88]. So far, the scientific literature has reported both terrestrial and aerial examples in which Robophysics played a pivotal role in enabling robots to adapt to real-world conditions by incorporating living-like mechanisms into their design.
@@
On the one side, on land, snake-like robots have been designed to move through unpredictable and confined spaces by mimicking the undulating motion of snakes, which use their body segments and frictional forces to generate propulsion [83, 84]. This ability makes snake-like robots highly suitable for deployment in disaster zones, where they can traverse rubble and debris to locate survivors or assess structural damage[74, 39]. Studies of snake locomotion across diverse terrains have allowed the creation of robots excelling in mobility, adaptability, and durability, even under extreme conditions [69, 86]. At a smaller length-scale, peristaltic worm-inspired soft robots have been designed for a better understanding of locomotion in different environments [29].
@@
On the other side, in the air, bird-inspired drones have been tailored to exploit flapping-wing aerodynamics to achieve maneuverability and energy efficiency which surpasses traditional fixed-wing or rotor designs[71]. At a smaller length scale, insect-inspired robots have been designed for crop pollination and environmental monitoring[24]. Such bio-inspired abstractions demonstrate how Robophysics, by bridging biology and engineering, enhances robotic adaptability in complex real-world environments [51].
@@
3 Learning from living systems: collective behaviour
@@
Living systems achieve collective functionality not via a central control, but via local interactions which generate robust and adaptable group behavior [47, 43]. Across length-scales, organisms actively consume energy to remain out of equilibrium and thereby dynamically respond to environmental cues [90]. Ranging from the micro to the macroscale, bacteria form dense swarms to withstand exposure to antibiotics [21], undergoing self-sustained active turbulence relevant for fluid mixing and molecular transport [33, 91, 80] , cytoskeletal filaments continuously assemble and disassemble to enable dynamic cell shape changes [37], while flocks of birds or schools of fish reorganize within fractions of a second to conserve energy or evade predators [63, 28, 72]. The unifying principle among all these systems is that coordinated responses emerge from simple, local rules of sensing and interaction, yet giving rise to collective resilience and adaptability.
@@
Efficient robot swarms should rely on fundamental processes-communication, coordination, cooperation, collaboration, and, at times, competition [73], to achieve collective goals. Algorithms based on these tasks have already achieved remarkable results in controlled experimental environments with a limited number of individuals in the swarm (fewer than ten). The following sections discuss each of these aspects and their relevance to collective robotic behavior.
3.1 Communication and its constraints
@@
Communication is the key for enabling swarm coordination. In robophysical collectives, it is fundamentally constrained by density, mobility, and morphology, which jointly determine the timeliness, reliability, and cost of information exchange. Gielis et al. [41] argue that communication in robot swarms confronts two major challenges. First, learning models should be employed not only to determine what information to communicate and to whom, but also optimize multiple aspects of swarm performance. Second, robots must be designed together with their communication capabilities, since any control algorithm relying on explicit communication is prone to failure when network data are insufficient and to sub-optimal performance when the network transmits excessive information.
Specifically, these constraints are regime-dependent: changes in density and morphology can abruptly alter contention, connectivity, and information freshness, thereby setting hard limits on coordination performance. For this reason, standard communication protocols cannot be adopted directly, because they aim at improving generic network performance metrics rather than ensuring the timeliness and scalability required by swarm coordination. Consequently, robophysical swarms call for task-driven, control-aware communication that explicitly limits transmissions and adapts what/when/to whom communicating (and at which rate) to maintain timeliness and scalability across varying density and morphology regimes.
More fundamentally, in Robophysics communication is not a passive substrate but part of the collective dynamics: information exchange affects motion and coordination, while motion continuously reshapes connectivity and contention. Designing mechanisms that explicitly account for this bidirectional coupling—and remain stable across regime shifts—remains an open challenge.
3.2 Coordination and Scalability
@@
In robotic swarms, objectives are collective and must be achieved through strictly local interactions with limited global knowledge, making coordination a central challenge. Coordination aims to minimize redundant actions and interference through cooperative behaviors such as trajectory coordination for area coverage and collision avoidance [94]. Scalability is essential as the number of robots increases, yet many techniques performing well for small groups become impractical at larger swarm sizes. Recent research has therefore focused on the formal control of shape formation in robot collectives [78]. In these systems, robots communicate with local neighbors to maintain a desired configuration, and their interactions are typically represented as a graph. In robophysical swarms, however, this graph is not merely a design abstraction but an outcome of morphology and density; as the collective transitions across regimes, graph connectivity and the available bandwidth per node can change abruptly, challenging the stability and convergence of formation and coverage controllers. Scaling up such formations requires enlarging the interaction graph while preserving key topological properties, a task that remains particularly challenging.
@@
FIG. 1: Schematic illustration of how principles from living systems can inform robophysics. In living systems (top), groups of organisms such as fish, ants, and birds display collective behavior, adaptability, and shared purpose. These principles inspire robotic collectives (bottom), where multi-robot swarms must address challenges such as communication, perception and sensing, and energy and resource management.
3.3 Networking and cooperation
At first glance, robophysical networking issues may appear similar to those in bio-inspired networking, as both draw on concepts from collective behavior in natural systems. However, bio-inspired networking primarily aims to leverage biological models to improve the performance of routing, coordination, synchronization and resource-allocation algorithms in large-scale communication networks [32]. Robophysics, instead, seeks to develop a physics-based understanding of how real robots move, interact, and generate collective behaviors under physical constraints.
From a computer networking perspective, the operation of a swarm of small, low-cost robots faces several fundamental challenges. Unlike abstract network nodes, robots are physical entities whose positions and motions directly affect network connectivity. As a result, the network topology continuously evolves with swarm motion, requiring tightly coupled strategies for locomotion, control, and communication. Crucially, many swarm networking functions (e.g., neighbor discovery, local coordination, and topology-aware communication) implicitly assume robots estimating the identity of their neighbors and the evolution of the interaction graph. Two notions are central in robophysical networking: embodiment (robots’physical motion and morphology shape connectivity, interference, and access to shared resources) and embeddedness(sensing, computation, and communication must be carried onboard). The latest happens under strict energy and form-factor constraints and often without external infrastructure (e.g., anchors, beacons, motion-capture, or centralized fusion). Importantly, these notions are coupled: embedding additional functionality (sensors, computation, radios, batteries) changes size, mass, and power draw, thereby affecting locomotion and, ultimately, the networked collective dynamics.
In location-deprived environments (such as indoor experiments) simultaneous localization and mapping (SLAM) becomes essential, Typical SLAM and localization pipelines often rely on assumptions which do not hold in robophysical swarms, such as the availability of external infrastructure (e.g., anchors), pre-installed maps, or stable reference frames. In fully infrastructure-free deployments, robots must instead infer relative positions and the interaction graph from sparse, noisy onboard sensing and intermittent communication. Performing topology mapping and cooperative localization in indoor environments with only simple sensors is thus a significant challenge. Relying on low-cost sensors inherently introduces sensor uncertainty, since measurements of distance, velocity, or orientation are often noisy or incomplete. This, in turn, leads to perception uncertainty, as obstacles or neighboring robots may be only partially detected or misidentified due to occlusions or reflections. Moreover, robot mobility exacerbates communication uncertainty, as time-varying and intermittent wireless links can cause packet losses and transmission delays.
@@ 3.3 Networking and cooperation
Cooperation refers to the ability of multiple (often homogeneous) agents to accomplish tasks that would be infeasible individually, such as jointly transporting a payload or assisting one another in traversing obstacles. For decades, computer networking has drawn inspiration from collective behavior in natural systems, leading to bio-inspired paradigms such as ant-colony-inspired routing and optimization schemes. In ant-based routing for mobile ad hoc networks, artificial “ants” (control or data packets) explore alternative paths and probabilistically reinforce successful ones, mimicking pheromone-mediated trail formation. This approach, known as Ant Colony Optimization (ACO), has proven effective in dynamic, decentralized settings by leveraging self-organization and adaptability to changing network topologies.
@@
Similarly, evolutionary algorithms such as genetic algorithms (GAs) have been widely used for path planning and task assignment in multi-robot and multi-drone systems. For instance, GA-based strategies such as GenPath [9] optimize UAV missions under constraints including limited battery life, intermittent communication, and multi-phase recharging, enabling scalable coordination in applications such as search-and-rescue, environmental monitoring, and delivery.
Despite their success, ACO and GA-based methods are formulated around offline or well-defined optimization objectives and do not fully capture the real-time, embodied responsiveness of biological collectives, where coordination emerges from local physical interactions and rapid adaptation under changing conditions and constraints. This gap motivates Robophysics as a complementary viewpoint: rather than imposing top-down coordination, it seeks to program an emergent collective behavior from the local interactions’ physics, fluctuations and energy consumption.
3.4 Collaboration, Learning and Adaptation
@@
Collaboration emphasizes teamwork leveraging the unique skills of different individuals, such as integrating various sensors or carrying out a common patrol task, combining aerial and aquatic autonomous vehicles.
@@
Learning concerns the swarm’s ability to learn from new scenarios and adjust its collective behaviour to achieve its goals, which is crucial for effective operation. While learning, robots in a swarm can collectively perceive their surroundings and share that information among themselves (distributed perception and sensing). In the paradigm of embodied evolution, for instance, robots continuously exchange and modify their own control “genomes” through local interactions, enabling real-time adaptation to the environment and to the collective dynamics themselves [18]. To better adapt to the environment, robots should exhibit both robustness and fault tolerance. Although these properties are related, they represent distinct concepts. Robustness refers to the ability to maintain satisfactory performance even when environmental conditions differ from those anticipated during the design phase. Fault tolerance, instead, refers to the capacity of the swarm to sustain effective operation despite failures in individual agents, sensors, or other components.
3.5 Competition and a shared purpose
@@
In both biological and robotic collectives, competition can coexist with cooperation, as individuals pursuing local goals may still contribute to global functionality. Such emergent coordination depends not only on feedback regulation but on a shared purpose. At the microscopic scale, gene regulatory networks couple local signals to coordinated multicellular responses during processes like morphogenesis, embodying distributed feedback [46, 6, 13]. At the macroscopic scale, local interaction rules (e.g., alignment, attraction, repulsion) are goal-conditioned: animals modulate the way they weigh neighbors and environmental cues in service of collective functions such as predator evasion, energy saving, or foraging [28, 72, 75, 82]. Robustness in living systems thus arises from distributed decision-making with a purpose, real-time adaptability driven by local feedback tuned to a common objective, rather than pre-planned, centralized control [64].
@@
3.6 Testing robot swarms